# Practical, Utilitarian Algorithm Configuration

### v1

@@ 3 COUP’s Empirical Performance
We have argued theoretically, and provide experimental evidence below, that our changes from Section 2 only improved COUP, first by tightening the bounds and refining the implementation of UCB (Figure 2) (Left) and second by incorporating a performance model (Figure 2) (Center). As mentioned in the introduction, even the previous version of COUP was shown by Graham and Leyton-Brown (2024) to broadly outperform other configuration procedures that offer theoretical guarantees; we refer the reader to Figure 5 of that work. We therefore did not perform an updated version of those (expensive) experiments. Instead, our aim in this paper is to compare COUP to the heuristic procedure that is most widely used in practice: SMAC. While SMAC was initially proposed for the optimization of expected runtime, it is able to optimize any objective function, and so can be straightforwardly extended to a utilitarian objective. We thus compared COUP’s performance to that of SMAC 3 (Lindauer et al. 2022) in scenarios consisting of four different SAT solvers and 20 different instance sets.
@@
We perform experiments using the ACLib library of algorithm configuration benchmarks (Hutter et al. 2014). ACLib is well-established as a benchmark set for demonstrating algorithm configuration performance. To the best of our knowledge, ours is the first paper to demonstrate algorithm configuration with general optimality guarantees in an online setting, using real runs from real solvers. In contrast, previous work has used precomputed or synthetic data.
@@
We ran experiments on a 14-node CPU cluster. Each node consisted of 32 dual-socket Intel Xeon E5-2683 CPUs with 96GB of memory. Each COUP scenario (solver and dataset combination) was trained for 96 hours with 12GB of memory per job, with each validation instance having a timeout of 60 seconds once training had been completed. We use the runsolver timing tool (Roussel 2011) to obtain runtime measurements that accurately reflected the total CPU time consumed by each solver.
@@
For each solver–instance set pair, we evaluated performance on a random sample of 2000 training instances (unless the dataset consisted of fewer than 2000 instances). After each iteration of COUP’s training, we saved the best configuration and validated each unique configuration found. We then evaluated SMAC’s incumbent configuration at the same points in time and on the same set of validation instances. We limited the individual solver runs that SMAC performed to 60 seconds and used the same limit when validating both COUP and SMAC. We used an initial captime of 1 second for COUP, which is the smallest execution time that runsolver is able to measure. We set COUP’s failure probability $\delta=0.01$.
@@
The utility function we used has two parameters: $\kappa_{0}$ and $\kappa_{1}$. If $t<\kappa_{0}$ then $u(t)=1$, indicating that all runs faster than $\kappa_{0}$ are identical to us and essentially instantaneous. If $t>\kappa_{1}$ then $u(t)=0$, indicating that all runs longer than $\kappa_{1}$ are essentially worthless to us. Between these two extremes, when $t\in[\kappa_{0},\kappa_{1}]$ we have $u(t)=\frac{\log(t/\kappa_{1})}{\log(\kappa_{0}/\kappa_{1})}$, meaning that if runtime is doubled then utility is reduced by a constant factor for any runtime in $\kappa_{0},\kappa_{1}]$. Optimizing this utility function, which amounts to a scaled and shifted logarithmic transform of runtime, will return configurations that are good for runtimes that are of different orders of magnitude, making it applicable over a broad range of problems. Moreover, using a logarithmic transform of runtime has been shown to improve the quality of algorithm runtime prediction models (Xu et al. 2008). We set $\kappa_{0}=0.001$ seconds, which is approximately the time it takes the operating system to launch a process and thus is the absolute minimum runtime we could observe. We set $\kappa_{1}=1$ hour, which is the maximum amount of time we imagine waiting for any one run to finish. Of course other utility functions are possible. We note that some utility functions might be more “strict” than others and will take more time to optimize. In general, the utility functions we consider drop from 1 to 0 as runtime increases. If this drop tends to happen at large runtime values, then COUP will tend to have to do longer runs, and take more time (depending on the given runtime distributions).
@@
Our augmented version of COUP identified good configurations about as quickly as SMAC at our cutoff time of four wall-clock days. For every (solver, instance) pair, we define the normalized utility as the utility achieved by a given algorithm configuration run divided by the best such utility observed across both the SMAC and COUP runs. Figure 2 (Right) is an empirical CDF of the difference between the normalized utilities achieved by COUP and SMAC across the different solver and instance set pairs. We call runs that achieved solutions within within 10% of each other effectively equivalent, represented as the region between the dashed red lines. $71$ out of our $80$ runs fell within this region; of what remains, $5$ had a difference less than 0.1-0.1 (meaning SMAC outperformed), and $4$ had a difference greater than $0.1$ (COUP outperformed). We performed the same analysis at various earlier cutoff times and observed qualitatively similar results.
@@
It is worth noting that COUP solves a considerably harder problem than SMAC. Rather than simply identifying strong algorithm configurations, it also proves anytime theoretical guarantees about the quality of the solution it has found at every point in time, both in terms of the amount by which the returned configuration might be worse than some other configuration under consideration ($\epsilon$) and in terms of the probability that a better configuration could be sampled ($\gamma$). Figure 3 shows the optimality guarantees proved by COUP for different solver–instance set pairs. Generally, both $\epsilon$ (Left) and $\gamma$ (Right) are driven toward 0, but some trends are evident. ProbSAT struggles to prove a good $\epsilon$ (purple curves tend to be higher). All solvers struggle to prove a good $\epsilon$ on at least one dataset (dark cluster at the top of the plot), and could use with more training time to push $\epsilon$ even farther down. Figure 3 (Right) begins with a solid line (hence the compressed $x$ axis) because COUP only tightens $\epsilon$ until it has learned sufficiently about its initial set of 100 configurations. Plots for individual runs can be found in Appendix C.
@@
Figure 3: Unlike SMAC, COUP also offers theoretical guarantees. In general, both $\epsilon$ and $\gamma$ are driven toward 0. Left: $\epsilon$ guarantee over time. Right: $\gamma$ guarantee over time.
@@
Finally, Graham and Leyton-Brown (2024) included a comparison of basic COUP to SuccessiveHalving (Jamieson and Talwalkar 2016), Hyperband (Li et al. 2018), AC-Band (Brandt et al. 2023), Structured Procrastination (Kleinberg et al. 2019), and ImpatientCapsAndRuns (Weisz et al. 2020). All of these procedures make theoretical guarantees of some sort, but it is important to note that these guarantees are not necessarily comparable to one another. COUP was found to perform well compared to all of these competitors, as was its finite-configuration-set variant, OUP. We included our improved variant of COUP (and OUP) in this comparison. Figure 7 in Appendix D plots the gap between the best configuration found and the best configuration overall for each of these procedures. It can be seen that our improved variants perform even better than their basic counterparts, in some cases quite significantly.
@@ 4 Utility Functions from Runtime CDFs
If the organizers decide that all runs less than 10 seconds are essentially equivalent, then this will be reflected in the chosen utility function, which will be constant up to (at least) 10 seconds. In this case we are essentially back to having FOSD. The winning solver is expected-utility preferred over solver 21 for any utility function that is constant up to 10 seconds. For instance one of the simplest evaluation criteria we could use in this scenario is the proportion of instances solved. This amounts to optimizing a step function utility of the form $u(t)=1$ if $t<\kappa$ and $u(t)=0$ otherwise, where $\kappa$ is a parameter. For any $\kappa>10$ seconds, the winning solver is FOSD over solver 21. In general, a utility function decreases from 1 to 0 as $t$ increases. The regions where the utility function decreases most will be the most important regions. If all of the decrease happens in a region where solver $A$ is FOSD over solver $B$, then $A$ will be expected utility preferred over $B$, regardless of the precise form of the decrease. We expand on this idea in Appendix B.
@@ 5 Conclusion
This paper has proposed four improvements to the utilitarian algorithm configuration procedure COUP that improve its empirical performance without degrading its theoretical bounds. We have shown that these improvements make a real difference to the tightness of the bounds proved by COUP, as well as to the quality of the returned configuration. We demonstrated evidence that COUP’s performance is competitive with SMAC, the most widely used heuristic configuration procedure, while also offering theoretical optimality guarantees that improve with time. We are now inclined to believe that utilitarian algorithm configuration methods that offer theoretical guarantees have reached the point that they should be preferred for use in practice. We also described a case study on the SAT Competition that illustrates how an algorithm evaluator can reason about the robustness of a conclusion recommended by a utilitarian configuration procedure with respect to different utility functions that could have been chosen, implying that uncertainty about the exact utility function used need not be a barrier to the adoption of utilitarian algorithm configuration approaches. In future work we plan to perform a more extensive experimental evaluation to make this claim definitive, comparing COUP’s performance with SMAC across a wider set of utility functions, and also contrasting COUP with other heuristic methods.
@@ Appendix A Algorithm Pseudocode
Pseudocode for the original COUP, as reported in Graham and Leyton-Brown (2024), is presented in Algorithm 1. Pseudocode for our improved version of COUP is presented in Algorithm 2. We discuss particular differences between the two with respect to the four improvements we have suggested. Define $d(p,q)=p\log(p/q)+(1-p)\log((1-p)/(1-q))$.
@@ A.1 Improved Confidence Bounds
The term $\alpha_{p}$ (defined in Graham and Leyton-Brown (2024)) is explicitly chosen to satisfy Hoeffding’s inequality in the bounds defined on lines 26 and 27 (and 12 and 13).
@@ Appendix B Deriving Confidence Bounds
Suppose we have upper- and lower- confidence bounds on the CDF $F$ at $\kappa$:
$F^{(lcb)}\leq F(\kappa)\leq F^{(ucb)}.$
It doesn’t matter where these bounds come from; they may be from Hoeffding’s inequality, or from the KL-term directly, or from somewhere else. Simple arithmetic then implies that
$1-F^{(lcb)}\geq 1-F(\kappa)\geq 1-F^{(ucb)}$
@@
$-u(\kappa)\big(1-F^{(lcb)}\big)\leq-u(\kappa)\big(1-F(\kappa)\big)\leq-u(\kappa)\big(1-F^{(ucb)}\big).$
@@
Using this, and the above bounds on $u(\kappa)\big(1-F(\kappa)\big)$, and the fact that $\mathbb{E}[u(t)|t>\kappa](1-F(\kappa))$ is positive, gives:
@@ Appendix D Comparison with Other Theoretically-motivated Procedures
Figure 7 shows a comparison of our improved variants COUP and OUP against a number of other theoretically-motivated configuration procedures. This is essentially a reproduction of Figure 5 from Graham and Leyton-Brown (2024) with out new COUP and OUP variants included.
@@
Figure 7: Figure 5 reproduced from Graham and Leyton-Brown (2024) with our refinements added to OUP and COUP. Top row: Procedures that can optimize any utility function (log-Laplace utility function was used). Bottom Row: Procedures that optimize some kind of capped average runtime.
@@ Appendix E Details of the Prediction Model
We use the XGBoost (Chen and Guestrin 2016) for implementing the prediction model. This allows us to predict a configuration’s mean utility. We produce uncertainty estimates by performing a bootstrap aggregation procedure. We subsample 100 sets of configurations of size twice the number of configurations, and train an XGBoost model on each one. For each candidate configuration we make a prediction with each of the 100 models, giving us a distribution of values. We then consider 95% confidence intervals, giving us an uncertainty band around each prediction. We do 100 rounds of boosting, giving 100 trees, which we limit to a depth of 3. We train the model with a squared error loss.

### v2

@@ 3 COUP’s Empirical Performance
We have argued theoretically, and provide experimental evidence below, that our changes from Section 2 only improved COUP, first by tightening the bounds and refining the implementation of UCB (Figure 2, Left) and second by incorporating a performance model (Figure 2, Center). We also compare COUP to the heuristic procedure that is most widely used in practice: SMAC. While SMAC was initially proposed for the optimization of expected runtime, it is able to optimize any objective function, and so can be straightforwardly extended to a utilitarian objective. We thus compared COUP’s performance to that of SMAC 3 (SMAC3) in scenarios consisting of four different SAT solvers and 20 different instance sets.
@@
We perform experiments using the ACLib library of algorithm configuration benchmarks (hutter2014aclib). ACLib is well-established as a benchmark set for demonstrating algorithm configuration performance. To the best of our knowledge, ours is the first paper to demonstrate utilitarian algorithm configuration with general optimality guarantees in an online setting, using real runs from real solvers. In contrast, previous work has used precomputed or synthetic data.
@@
We ran experiments on a 14-node CPU cluster. Each node consisted of 32 dual-socket Intel Xeon E5-2683 CPUs with 96GB of memory. Each COUP scenario (solver and dataset combination) was trained for 96 hours with 12GB of memory per job, with each validation instance having a timeout of 60 seconds once training had been completed. We use the runsolver timing tool (roussel2011controlling) to obtain runtime measurements that accurately reflected the total CPU time consumed by each solver.
@@
For each solver–instance set pair, we evaluated performance on a random sample of 2000 training instances (unless the dataset consisted of fewer than 2000 instances). After each iteration of COUP’s training, we saved the best configuration and validated each unique configuration found. We then evaluated SMAC’s incumbent configuration at the same points in time and on the same set of validation instances. We limited the individual solver runs that SMAC performed to 60 seconds and used the same limit when validating both COUP and SMAC. We used an initial captime of 1 second for COUP, which is the smallest execution time that runsolver is able to enforce. We set COUP’s failure probability $\delta=0.01$.
@@
The utility function we used has two parameters: $\kappa_{0}$ and $\kappa_{1}$. If $t<\kappa_{0}$ then $u(t)=1$, indicating that all runs faster than $\kappa_{0}$ are identical to us and essentially instantaneous. If $t>\kappa_{1}$ then $u(t)=0$, indicating that all runs longer than $\kappa_{1}$ are essentially worthless to us. Between these two extremes, when $t\in[\kappa_{0},\kappa_{1}]$ we have $u(t)=\frac{\log(t/\kappa_{1})}{\log(\kappa_{0}/\kappa_{1})}$, meaning that if runtime is doubled then utility is reduced by a constant factor for any runtime in $\kappa_{0},\kappa_{1}]$. Optimizing this utility function, which amounts to a scaled and shifted logarithmic transform of runtime, will return configurations that are good for runtimes that are of different orders of magnitude, making it applicable over a broad range of problems. Moreover, using a logarithmic transform of runtime has been shown to improve the quality of algorithm runtime prediction models (xu2008satzilla). We set $\kappa_{0}=0.001$ seconds, which is approximately the time it takes the operating system to launch a process and thus is the absolute minimum runtime we could observe. We set $\kappa_{1}=1$ hour, which is the maximum amount of time we imagine waiting for any one run to finish. Of course other utility functions are possible. We note that some utility functions might be more “strict” than others and will take more time to optimize. In general, the utility functions we consider drop from 1 to 0 as runtime increases. If this drop tends to happen at large runtime values, then COUP will tend to have to do longer runs, and take more time (depending on the given runtime distributions).
@@
Our augmented version of COUP identified good configurations about as quickly as SMAC at our cutoff time of four wall-clock days. For every (solver, instance) pair, we define the normalized utility as the utility achieved by a given algorithm configuration run divided by the best such utility observed across both the SMAC and COUP runs. Figure 2 (Right) is an empirical CDF of the difference between the normalized utilities achieved by COUP and SMAC across the different solver and instance set pairs. We call runs that achieved solutions within 10% of each other effectively equivalent, represented as the region between the dashed red lines. $71$ out of our $80$ runs fell within this region; of what remains, $5$ had a difference less than 0.1-0.1 (meaning SMAC outperformed), and $4$ had a difference greater than $0.1$ (COUP outperformed). We performed the same analysis at various earlier cutoff times and observed qualitatively similar results.
@@
It is worth noting that COUP solves a considerably harder problem than SMAC. Rather than simply identifying strong algorithm configurations, it also proves anytime theoretical guarantees about the quality of the solution it has found at every point in time, both in terms of the amount by which the returned configuration might be worse than some other configuration under consideration ($\epsilon$) and in terms of the probability that a better configuration could be sampled ($\gamma$). Figure 3 shows the optimality guarantees proved by COUP for different solver–instance set pairs. Generally, both $\epsilon$ (Left) and $\gamma$ (Right) are driven toward 0, but some trends are evident. ProbSAT struggles to prove a good $\epsilon$ (purple curves tend to be higher). All solvers struggle to prove a good $\epsilon$ on at least one dataset (dark cluster at the top of the plot), and could use with more training time to push $\epsilon$ even farther down. Figure 3 (Right) begins with a solid line (hence the compressed $x$ axis) because COUP tightens only $\epsilon$ until it has learned sufficiently about its initial set of 100 configurations. Plots for individual runs can be found in Appendix G.
@@
Figure 3: Unlike SMAC, COUP also offers theoretical guarantees. In general, both $\epsilon$ and $\gamma$ are driven toward 0.
@@
Finally, graham2024utilitarian included a comparison of basic COUP to SuccessiveHalving (jamieson2016non), Hyperband (li2018hyperband), AC-Band (brandt2023ac), Structured Procrastination (kleinberg2019procrastinating), and ImpatientCapsAndRuns (weisz2020impatientcapsandruns). All of these procedures make theoretical guarantees of some sort, but it is important to note that these guarantees are not necessarily comparable to one another. COUP was found to perform well compared to all of these competitors, as was its finite-configuration-set variant, OUP. We included our improved variant of COUP (and OUP) in this comparison. Figure 7 in Appendix D plots the gap between the best configuration found and the best configuration overall for each of these procedures. It can be seen that our improved variants perform even better than their basic counterparts, in some cases quite significantly.
@@ 4 Utility Functions from Runtime CDFs
If the organizers decide that all runs less than 10 seconds are essentially equivalent, then this will be reflected in the chosen utility function, which will be constant up to (at least) 10 seconds. In this case we are essentially back to having FOSD. The winning solver is expected-utility preferred over solver 21 for any utility function that is constant up to 10 seconds. For instance one of the simplest evaluation criteria we could use in this scenario is the proportion of instances solved. This amounts to optimizing a step function utility of the form $u(t)=1$ if $t<\kappa$ and $u(t)=0$ otherwise, where $\kappa$ is a parameter. For any $\kappa>10$ seconds, the winning solver is FOSD over solver 21. In general, a utility function decreases from 1 to 0 as $t$ increases. The regions where the utility function decreases most will be the most important regions. If all of the decrease happens in a region where solver $A$ is FOSD over solver $B$, then $A$ will be expected utility preferred over $B$, regardless of the precise form of the decrease. We expand on this idea in Section F.1.
@@ 5 Conclusion
We have proposed four improvements to the utilitarian algorithm configuration procedure COUP that boost its empirical performance without degrading its theoretical bounds. We have shown that these improvements make a real difference to the tightness of the bounds proved by COUP, as well as to the quality of the returned configuration. We demonstrated that COUP is competitive with SMAC, the most widely used heuristic configuration procedure, while also offering theoretical optimality guarantees that improve with time. We are now inclined to believe that utilitarian algorithm configuration methods that offer theoretical guarantees have reached the point that they should be preferred for use in practice. We also described a case study on the SAT Competition that shows how to reason about the robustness of a conclusion recommended by a utilitarian configuration procedure with respect to different utility functions that could have been chosen, implying that uncertainty about the exact utility function used need not be a barrier to the adoption of utilitarian algorithm configuration approaches. In future work we plan to perform a more extensive experimental evaluation to make this claim definitive, comparing COUP’s performance with SMAC across a wider set of utility functions, and also contrasting COUP with other heuristic methods.
@@ Appendix A Algorithm Pseudocode
Pseudocode for the original COUP, as reported in graham2024utilitarian, is presented in Algorithm 1. Pseudocode for our improved version of COUP is presented in Algorithm 2. We discuss particular differences between the two with respect to the four improvements we have suggested. Define $d(p,q)=p\log(p/q)+(1-p)\log((1-p)/(1-q))$.
@@ A.1 Improved Confidence Bounds
The term $\alpha_{p}$ (defined in graham2024utilitarian) is explicitly chosen to satisfy Hoeffding’s inequality in the bounds defined on lines 26 and 27 (and 12 and 13).
@@ Appendix B Deriving Confidence Bounds
Suppose we have a lower confidence bounds on the CDF $F$ at $\kappa$:
$F^{(lcb)}\leq F(\kappa).$
It doesn’t matter where this bound comes from; it may be from Hoeffding’s inequality, or from the KL-term directly, or from somewhere else. Simple arithmetic then implies that
$1-F^{(lcb)}\geq 1-F(\kappa)$
@@
$-u(\kappa)\big(1-F^{(lcb)}\big)\leq-u(\kappa)\big(1-F(\kappa)\big).$
@@
Using this, and the above bound on $u(\kappa)\big(1-F(\kappa)\big)$, and the fact that $\mathbb{E}[u(t)|t>\kappa](1-F(\kappa))$ is positive, gives:
@@ Appendix D Comparison with Other Theoretically-motivated Procedures
Figure 7 shows a comparison of our improved variants COUP and OUP against a number of other theoretically-motivated configuration procedures. This is essentially a reproduction of Figure 5 from graham2024utilitarian with out new COUP and OUP variants included.
@@
Figure 7: Figure 5 reproduced from graham2024utilitarian with our refinements added to OUP and COUP. Top row: Procedures that can optimize any utility function (log-Laplace utility function was used). Bottom Row: Procedures that optimize some kind of capped average runtime.
@@ Appendix E Details of the Prediction Model
We use the XGBoost (chen2016xgboost) for implementing the prediction model. This allows us to predict a configuration’s mean utility. We produce uncertainty estimates by performing a bootstrap aggregation procedure. We subsample 100 sets of configurations of size twice the number of configurations, and train an XGBoost model on each one. For each candidate configuration we make a prediction with each of the 100 models, giving us a distribution of values. We then consider 95% confidence intervals, giving us an uncertainty band around each prediction. We do 100 rounds of boosting, giving 100 trees, which we limit to a depth of 3. We train the model with a squared error loss.